OpenStudy (anonymous):

Zeus Industries bought a computer for $2859. It is expected to depreciate at a rate of 21% per year. What will the value of the computer be in 3 years?

OpenStudy (anonymous):

@Nnesha think you could help me with this one?

Nnesha (nnesha):

depreciate means decrease so you have to use this formula \[\huge\rm y = a(1-r)^t\] a = initial amount r= rate t = years substitute values solve for y

Nnesha (nnesha):

and first you need to divide 21/100 change percent to decimal

OpenStudy (anonymous):

0.21

Nnesha (nnesha):

yes right so r =.21

OpenStudy (anonymous):

so the equation looks like y=2859(1-.21)^3

Nnesha (nnesha):

ye sright

OpenStudy (anonymous):

so now just solve it?

Nnesha (nnesha):

yes

OpenStudy (anonymous):

1409.59?

Nnesha (nnesha):

looks right

OpenStudy (anonymous):

and rounded to the nearest cent would be 1409.60? if it was1409.598?

OpenStudy (anonymous):

could you help me with 4 more please?

OpenStudy (anonymous):

A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a/2=ae^−3.568t

Nnesha (nnesha):

i have to go close this post and make new one so other people able to help u sorry